OpenStudy (anonymous):

Brandy is making a peach pie for her grandmother for Christmas. The diameter of the pie is 12 in. If three-fourths of the area of the pie was eaten after Christmas dinner, what is the area of the pie that is leftover?

OpenStudy (anonymous):

the area of the pie before it was eaten = pi * 6^2 in^2

OpenStudy (mathmagician):

A=(1-3/4)*(3.14*6^2)=28.26 sq. in.

OpenStudy (anonymous):

THANKS AGAIN